the sum of an integer and its square is 72. What's the integer?
Answer by oberobic(2304)

x +x^2 = 72
.
x^2 +x -72 = 0
.
factor
.
(x+9)(x-8) = 0
.
x = -9 or 8
.
check both values.
.
-9^2 = 81
81+(-9) = 72
correct
.
8^2 = 64
64 + 8 = 72
correct
.
The integer is either -9 or 8.
.
Done.